Huggins Co. has identified an investment project with the following cash flows.

YEAR Cash Flow
1 $740
2 $970
3 $1230
4 $1325

If the discount rate is 9 percent, what is the present value of these cash flows? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

Present value $ What is the present value at 18 percent? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

Present value $ What is the present value at 24 percent? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

If the discount rate is 9 percent, what is the present value of these cash flows

=740/(1+9%)^1+970/(1+9%)^2+1230/(1+9%)^3+1325/(1+9%)^4

=3383.78

If the discount rate is 18 percent, what is the present value of these cash flows

=740/(1+18%)^1+970/(1+18%)^2+1230/(1+18%)^3+1325/(1+18%)^4

=2755.79

If the discount rate is 24 percent, what is the present value of these cash flows

=740/(1+24%)^1+970/(1+24%)^2+1230/(1+24%)^3+1325/(1+24%)^4

=2433.19

the above is answer..